Solve.
The length of a rectangular room is 4 feet longer than twice the width. If the room's perimeter is 140 feet, what are the room's dimensions?

Answer by doukungfoo(195)

The perimeter of a rectangle is equal to the sum of all its sides.
A rectangle has two sides that represent its length and two sides that represent its width.
So with this knowledge about a rectangle we can write a formula:
2(Length) + 2(Width) = Perimeter
Thats 2 times the length plus 2 times the widthe equals perimeter.
Now we are told that the length of a rectangular room is 4 ft longer than twice its width.
Lets set width equal to x.
Width = x
Now define the length in terms of x
Length = 2x + 4
We are given the perimeter is 140 feet.
Now we can plug all this information into our formula and solve for x:
2(Length) + 2(Width) = Perimeter
2(2x+4) + 2(x) = 140
4x + 8 + 2x = 140
6x + 8 = 140
6x = 132
x = 22
Answer: the width is 22 ft
Now find the length
Length = 2x + 4
Length = 2(22) + 4
Length = 44 + 4
Length = 48
Answer: the length is 48 ft
Check answers:
2(Length) + 2(Width) = Perimeter
2(48) + 2(22) = 140
96 + 44 =140
140 = 140